English

If the probability of winning a game is 0.07, what is the probability of losing it?

Advertisements
Advertisements

Question

If the probability of winning a game is 0.07, what is the probability of losing it?

Sum
Advertisements

Solution

P(winning the game) = 0.07

P(losing the game) = 1 – 0.07

= 0.93

shaalaa.com
  Is there an error in this question or solution?
Chapter 16: Probability - VERY SHORT ANSWER TYPE QUESTIONS (VSAQS) [Page 16.29]

APPEARS IN

R.D. Sharma Mathematics [English] Class 10
Chapter 16 Probability
VERY SHORT ANSWER TYPE QUESTIONS (VSAQS) | Q 19. | Page 16.29
Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×